OpenStudy (lollola):

What are the values of the function y=3x -4 for x= 0,1,2 and 3?

OpenStudy (anonymous):

Okay, I'll help you out. All we have to do is substitute the exponent for one of the given values of x. By the way, is that 3x - 4 or 3x times negative 4?

OpenStudy (lollola):

- 4

OpenStudy (anonymous):

Okay great. Let's substitute x for 0. 3 x 0 = 0. 0 - 4 = -4. Now we understand that y = -4. Next let's do 1. 3 x 1 = 3. 3 - 4 = -1. y must equal -1. Can you solve the next two?

OpenStudy (lollola):

Im not really understanding still

OpenStudy (anonymous):

Oh, okay. I'll try to explain it. When we have a problem with an value, like x, that's just an unknown value. Like 1 + x = 1. We don't know what x is, and in this case, we won't know. But, we can substitute x for the numbers that were given, 0 1 2 3. Then we just treat it like a normal equation following the order of operations. Does it make any more sense?

OpenStudy (lollola):

A bit, yes

OpenStudy (lollola):

OH im starting to get it now!

OpenStudy (anonymous):

Okay, then I'll go through step by step solving the first problem. We have the problem \[y = 3x - 4\] Let's ignore the x and replace it with a 0. 3 x 0 = 0. Now, instead of \[y = 3x - 4\] We are left with this \[y = 0 - 4\] We can then just subtract 0 - 4 and get -4. Now we are finally left with this \[y = -4\] So, y must equal -4.

OpenStudy (lollola):

I get what your saying, but not sure how to answer the question still xP

OpenStudy (lollola):

Would I put the answers to the equations we are making with the substitutions for x?

OpenStudy (anonymous):

Hm... Well, the answer to the first equation would be y = -4, so I guess so.

OpenStudy (lollola):

Ok thanks for the help :)

OpenStudy (anonymous):

Of course! If you have any more questions, feel free to message me. And remember, if I helped you in any way, a fan and medal would be appreciated. :)

OpenStudy (lollola):

Of course!